# Cosmological Stasis from a Single Annihilating Particle Species: Extending Stasis Into the Thermal Domain

### v1

@@ VIII A Particle-Physics Mechanism for Achieving Thermal Stasis
The cosmological implications of the criterion in Eq. (50) for stasis in can be understood as follows. Let us assume, as we did in Sect. II, that additional scattering processes in the theory serve to maintain kinetic equilibrium among the population of $\phi$ particles, and that this population of particles can be characterized by a temperature $T$. At values of $T$ for which the phase-space integrals in $\langle\sigma v\rangle$ and $\langle({\rm KE}_{a}+{\rm KE}_{b})\sigma v\rangle$ are each dominated by the contribution from particles with $|\vec{p}_{\rm CM}|$ within the range specified in Eq. (53), these thermal averages are well approximated by the expressions in Eq. (24) with $q=-2$. Thus, at such temperatures, the universe evolves toward stasis. By contrast, at values of $T$ for which these phase-space integrals receive sizable contributions from particles with $|\vec{p}_{\rm CM}|$ outside this range, $\langle\sigma v\rangle$ and $\langle({\rm KE}_{a}+{\rm KE}_{b})\sigma v\rangle$ do not exhibit the appropriate scaling behavior with $T$ and the system does not evolve toward stasis.
@@ IX Conclusions and Directions for Future Research
In all of its realizations, cosmological stasis is ultimately a consequence of processes which transfer energy density from one cosmological component to another in a way which counteracts the effects of Hubble expansion. A variety of processes, including particle decay Dienes et al. (2022a, 2023), Hawking radiation Barrow et al. (1991); Dienes et al. (2022b), and the overdamped/underdamped transition of homogeneous scalar-field zero-modes Dienes et al. (2024) give rise to pumps which can accomplish this feat. However, the pump rate associated with each of these processes depends only on the intrinsic properties (e.g., masses or decay widths) of individual objects — individual particles, black holes, etc. — within the system, or on the dynamical variables $\Omega_{i}$ and $H$, or on $t$ itself. In other words, the pump rate was independent of the external environment in which these constituent objects found themselves.
@@
Second, despite the absence of a tower of states, the realization of thermal stasis that we have described in Sect. VIII comes with its own graceful exit. In tower-based realizations of stasis, the stasis epoch stretches across a time interval during which the action of the pump incrementally proceeds towards the lower portions of the tower Dienes et al. (2022a, 2023, 2024). The exit from the stasis epoch then arises once we reach the bottom of the tower. As a result, once stasis ends, the universe subsequently becomes dominated by that cosmological component which participates in the stasis dynamics and which has the highest equation-of-state parameter. For thermal stasis, by contrast, there is no tower. We nevertheless continue to have an extended stasis epoch, and this in turn ends through a different mechanism: once the temperature of our $\phi$-particle gas eventually drops below the temperature $T_{\rm min}$ in Eq. (55), $\langle\sigma v\rangle$ no longer increases as the kinetic-energy density of the $\phi$ particles decreases. As a result, $\phi\phi\to\chi\chi$ annihilation becomes inefficient and $\Omega_{M}$ increases until the $\phi$-particle gas (i.e., the component with the lowest equation-of-state parameter) ultimately comes to dominate the energy density of the universe. Thus, in this scenario, stasis ends not with a bang, but with a WIMP-er. Clearly the graceful exits for thermal and tower-based realizations of stasis are qualitatively quite different. We nevertheless see that each of our stasis epochs comes with its own intrinsic exit.

### v2

@@ VIII A Particle-Physics Mechanism for Achieving Thermal Stasis
The cosmological implications of the criterion in Eq. (50) for stasis can be understood as follows. Let us assume, as we did in Sect. II, that additional scattering processes in the theory serve to maintain kinetic equilibrium among the population of $\phi$ particles, and that this population of particles can be characterized by a temperature $T$. At values of $T$ for which the phase-space integrals in $\langle\sigma v\rangle$ and $\langle({\rm KE}_{a}+{\rm KE}_{b})\sigma v\rangle$ are each dominated by the contribution from particles with $|\vec{p}_{\rm CM}|$ within the range specified in Eq. (53), these thermal averages are well approximated by the expressions in Eq. (24) with $q=-2$. Thus, at such temperatures, the universe evolves toward stasis. By contrast, at values of $T$ for which these phase-space integrals receive sizable contributions from particles with $|\vec{p}_{\rm CM}|$ outside this range, $\langle\sigma v\rangle$ and $\langle({\rm KE}_{a}+{\rm KE}_{b})\sigma v\rangle$ do not exhibit the appropriate scaling behavior with $T$ and the system does not evolve toward stasis.
@@
Of course, any particle-physics model which gives rise to an annihilation process of the form illustrated in Fig. 6 necessarily also gives rise to a number of scattering processes which can in principle impact the cosmological dynamics of our system. Two such processes are $\phi\phi\to\phi\phi$ scattering and $\chi\chi\to\chi\chi$ scattering, which, if efficient, can serve to establish and maintain kinetic equilibrium among the cosmological populations of $\phi$ and $\chi$ particles, respectively. However, it turns out that within the particle-physics models that are capable of achieving thermal stasis via the mechanism described above, $\phi\phi\to\phi\phi$ scattering is typically efficient whereas $\chi\chi\to\chi\chi$ scattering typically is not. This is because this mechanism can only give rise to a stasis epoch with a significant duration when $T_{\rm min}\ll T_{\rm max}$ and the condition in Eq. (54) is therefore satisfied across a broad range of $T$. This typically requires that the $\phi$ particles couple to the mediator $X$ with a far greater strength than do the $\chi$ particles Barber et al. (2024). As a result, the population of $\phi$ particles remains in kinetic equilibrium in stasis models of this sort, whereas the population of $\chi$ particles produced by $\phi\phi\to\chi\chi$ annihilation typically never attains kinetic equilibrium and therefore possesses a highly non-thermal phase-space distribution.
A third process which necessarily arises in particle-physics models of this sort is $\chi\phi\to\chi\phi$ scattering — a process which facilitates the transfer of kinetic energy from radiation to matter. Scattering processes of this sort can potentially have a significant impact on the manner in which the temperature of a cosmological population of non-relativistic particles evolves over time in scenarios with non-standard expansion histories. For example, the manner in which a population of cold dark-matter particles kinetically decouples from the radiation bath during an early matter-dominated era (EMDE) can be very different from the manner in which such a population of particles decouples during a radiation-dominated era. In particular, due to the continuous injection of radiation from the heavy decaying particles which dominate the energy density of the universe during the EMDE, this decoupling process can include a lengthy quasi-decoupling phase wherein the temperature of the matter particles evolves in a non-standard manner Waldstein et al. (2017). One might therefore worry that a process like $\chi\phi\to\chi\phi$ scattering could alter the manner in which $T$ evolves with time in our thermal-stasis scenario such that the system does not in fact gives rise to stasis at all.
The impact that $\chi\phi\to\chi\phi$ scattering has on the evolution of $T$ within any particular particle-physics model in which our mechanism for achieving thermal stasis is realized depends sensitively on the details of the model. In what follows, we summarize the qualitative manner in which this process affects the cosmological dynamics. A more detailed, quantitative analysis, performed within the context of a minimal such model, is provided in Ref. Barber et al. (2024). It turns out that $\chi\phi\to\chi\phi$ scattering modifies the evolution of $T$ in thermal-stasis scenarios of this sort by giving rise to an $\Omega_{M}$-dependent correction to the coefficient of the first term in Eq. (11). In particular, in the presence of such scattering, this equation is modified to
[…]
$~{}~{}~{}~{}~{}-\,\frac{2m}{3\Omega_{M}}\left(P_{{\rm KE},\gamma}-\frac{\Omega_{\rm KE}}{\Omega_{M}}\,P_{M,\gamma}\right)~{},~{}~{}~{}~{}~{}~{}$
where $\epsilon$ is a constant factor which depends on the couplings involved. This correction term leads in turn to a correction term in the dynamical equation for $\Xi$ in Eq. (28).
Since $\chi\phi\to\chi\phi$ scattering proceeds through a $t$-channel process and therefore does not receive the sizable resonant enhancement that $\phi\phi\to\chi\chi$ annihilation receives, it is often the case that $\epsilon$ is extremely small. Indeed, one finds that the correction term in Eq. (56) is often negligible within the parameter-space regions of interest in thermal-stasis models of this sort and can therefore be ignored. Interestingly, however, it also turns out that the structure of this correction term is such that stasis emerges in this system and remains a global dynamical attractor even when $\epsilon$ is sufficiently large that this term cannot be neglected. Indeed, this term simply leads to a modification of the resulting stasis values ${\overline{\Omega}}_{M}$ and ${\overline{\Xi}}$. Thus, while $\chi\phi\to\chi\phi$ scattering enriches the cosmological dynamics in such model realizations of thermal stasis, it does not impact our qualitative results.
@@ IX Conclusions and Directions for Future Research
In all of its realizations, cosmological stasis is ultimately a consequence of processes which transfer energy density from one cosmological component to another in a way which counteracts the effects of Hubble expansion. A variety of processes, including particle decay Dienes et al. (2022a, 2024a), Hawking radiation Barrow et al. (1991); Dienes et al. (2022b), and the overdamped/underdamped transition of homogeneous scalar-field zero-modes Dienes et al. (2024b) give rise to pumps which can accomplish this feat. However, the pump rate associated with each of these processes depends only on the intrinsic properties (e.g., masses or decay widths) of individual objects — individual particles, black holes, etc. — within the system, or on the dynamical variables $\Omega_{i}$ and $H$, or on $t$ itself. In other words, the pump rate was independent of the external environment in which these constituent objects found themselves.
@@
Second, despite the absence of a tower of states, the realization of thermal stasis that we have described in Sect. VIII comes with its own graceful exit. In tower-based realizations of stasis, the stasis epoch stretches across a time interval during which the action of the pump incrementally proceeds towards the lower portions of the tower Dienes et al. (2022a, 2024a, 2024b). The exit from the stasis epoch then arises once we reach the bottom of the tower. As a result, once stasis ends, the universe subsequently becomes dominated by that cosmological component which participates in the stasis dynamics and which has the highest equation-of-state parameter. For thermal stasis, by contrast, there is no tower. We nevertheless continue to have an extended stasis epoch, and this in turn ends through a different mechanism: once the temperature of our $\phi$-particle gas eventually drops below the temperature $T_{\rm min}$ in Eq. (55), $\langle\sigma v\rangle$ no longer increases as the kinetic-energy density of the $\phi$ particles decreases. As a result, $\phi\phi\to\chi\chi$ annihilation becomes inefficient and $\Omega_{M}$ increases until the $\phi$-particle gas (i.e., the component with the lowest equation-of-state parameter) ultimately comes to dominate the energy density of the universe. Thus, in this scenario, stasis ends not with a bang, but with a WIMP-er. Clearly the graceful exits for thermal and tower-based realizations of stasis are qualitatively quite different. We nevertheless see that each of our stasis epochs comes with its own intrinsic exit.